# Learning Decision-Focused Uncertainty Sets in Robust Optimization

### v4

@@ Appendix D Proof of convergence of the augmented Lagrangian algorithm
From step 1, we have a stationary point $(\alpha^{k},s^{k},\theta^{k})$ of the augmented Lagrangian function at iteration $k$ of Algorithm 1, with fixed multiplier and penalty parameters $\lambda^{k-1}$ and $\mu^{k-1}$. Further properties of this stationary point depend on the convergence, or lack thereof, of the sequence $\{\lambda^{k}\}$. As the step size parameter of $\lambda$ is the penalty $\mu$, we consider the two cases, (a) where the sequence of penalty parameters $\{\mu^{k}\}$ is bounded, and (b) where $\{\mu^{k}\}$ is unbounded. We first analyze case (a). By lines 6-12 of Algorithm 1, we note that for $\{\mu^{k}\}$ to be bounded, the penalty parameter must remain constant from some iteration onwards. By the condition on line 6, with $\tau<1$, this implies $\lim_{k\rightarrow\infty}\|H(\alpha^{k},\theta^{k})+s^{k}-\kappa\|_{2}=0$, which in turn implies $\lim_{k\rightarrow\infty}\mu^{k}(H(\alpha^{k},\theta^{k})+s^{k}-\kappa)=0$. This sequence of updates for $\lambda$ is geometric and summable, so the sequence $\{\lambda^{k}\}$ is bounded, with a convergent subsequence. Given the boundedness assumption on the true Lagrangian multiplier $\lambda^{\star}$ in Assumption 3.4, we conclude that the whole sequence converges, ${\it i.e.},$$\{\lambda^{k}\}\rightarrow\lambda^{\star}$. We now turn to case (b), where $\{\mu^{k}\}$ is unbounded. In this case, $\{\lambda^{k}\}$ may not converge. However, by the $\lambda$ update rule in lines 7 of Algorithm 1, we note that while $\{\mu^{k}\}\rightarrow\infty$, $\{\lambda^{k}\}$ is bounded between $\lambda^{\rm{min}}$ and $\lambda^{\rm{max}}$, the safeguard values.
@@
We now show that either the sequence $\{\theta^{k}\}$ converges to a feasible point of the $\widehat{\mathop{\bf CVaR}}$ condition, or $\{\alpha^{k},s^{k},\theta^{k}\}$ converges to a stationary point of the constraint violation function. We begin with the former. Suppose the sequence of penalty parameters $\{\mu^{k}\}$ is bounded, and converges to $\mu^{\star}$. Then, as $k\to\infty$, $\{\lambda^{k}\}$ converges to the stationary point $\lambda^{\star}$. From step 1, we have that $\{\theta^{t}\}$ in the inner loop converges almost surely to $\theta^{k}=\theta^{\star}(\lambda^{k-1},\mu^{k-1})$ along each step of $\lambda$ and $\mu$’s convergence. In order to establish the convergence of $\{\theta^{k}\}$ in the outer loop to a stationary point $\theta^{\star}(\lambda^{\star},\mu^{\star})$, a feasible point of the $\widehat{\mathop{\bf CVaR}}$ condition (9), we must ensure that $\theta^{\star}(\lambda,\mu)$ is well-behaved, by showing its continuity with respect to $(\lambda,\mu)$.
@@
Recall Assumption 3.3 on the uniqueness of the inner RO solutions $x(\theta,y)$, Assumption 3.4 on the compactness of the set $\mathcal{X}$, and the convexity (and therefore continuity) of the objective and constraint functions $f$ and $g$ in $x$. Suppose these assumptions hold. By Theorem 4.2.1 of Bank et al., (1984), we establish the continuity of $L(\alpha,s,\theta,\lambda,\mu)$ in $\theta$, for $\theta\in\mathcal{D}_{\theta}$. As $L$ is bilinear and therefore continuous in $(\lambda,\mu)$, we obtain our desired stability results, i.e., as $\{\lambda^{k}\}$ and $\{\mu^{k}\}$ converge to $\lambda^{\star}$, $\mu^{\star}$, the optimal solution with respect to $\theta$ is continuous, and converges almost surely to the local minimizer $\theta^{\star}$. In addition, by the strong convexity of $L$ in $s$, the optimal solution with respect to $s$ also converges to the global minimizer $s^{\star}$. Note that $\{\alpha^{k}\}$ does not necessarily converge to a stationary point, as the minimum over $\alpha$ is not necessarily unique (Uryasev and Rockafellar,, 2001, Theorem 2). However, by the convergence of the inner loop in step 1 and by the convergence condition of $\mu$ in step 2, as $k\to\infty$ in Algorithm 1, we have a point $(\alpha^{k},s^{\star},\theta^{\star},\lambda^{\star},\mu^{\star})$ satisfying $0\in J_{L}(\alpha^{k},s^{\star},\theta^{\star})$ and $H(\alpha^{k},\theta^{\star})+s^{\star}-\kappa=0$. We can conclude that $\{\theta^{k}\}$ converges almost surely to a feasible point $\theta^{\star}$ of the $\widehat{\mathop{\bf CVaR}}$ condition, where the minimizer over $\alpha$ takes some value $\alpha^{\star}\leq\alpha^{k}$.
@@
On the other hand, suppose the sequence of penalty parameters $\{\mu^{k}\}$ is unbounded. By step 1, for any $\lambda^{k}$ and $\mu^{k}$, we have obtained a stationary point of the augmented Lagrangian function, satisfying $0\in J_{L}(\alpha^{k},s^{k},\theta^{k})$. Note that, with $\{\mu^{k}\}$ unbounded and $\{\lambda^{k}\}$ bounded, we observe $\lim_{k\rightarrow\infty}\lambda^{k}/\mu^{k}=0$. Therefore, dividing both sides of the conservative Jacobian by $\mu^{k}$ and taking the limit, we obtain:
@@
00 0 $\in\lim_{k\rightarrow\infty}\frac{1}{\eta NJ}\sum_{y^{j}\in Y^{J}}\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha^{k}\right\}((g(x^{j},u^{i},y^{j})-\alpha^{k})_{+}/\eta+\alpha^{k}+s^{k}-\kappa)J_{g}(x^{j})J_{x^{j}}(\theta^{k}),$
00 0 $\in\lim_{k\rightarrow\infty}\frac{1}{NJ}\sum_{(u^{i},y^{j})\in(U^{N}\times Y^{J})}(g(x^{j},u^{i},y^{j})-\alpha^{k})_{+}/\eta+\alpha^{k}+s^{k}-\kappa,$
@@
00 0 $\in\lim_{k\rightarrow\infty}\frac{1}{NJ}\sum_{(u^{i},y^{j})\in(U^{N}\times Y^{J})}\left(1-(1/\eta)\mathbf{1}\left\{g(x^{j},u^{i},y^{j})\geq\alpha^{k}\right\}\right)((g(x^{j},u^{i},y^{j})-\alpha^{k})_{+}/\eta+\alpha^{k}+s^{k}-\kappa).$
@@
These are the stationary conditions for the constraint violation function $V(\alpha,s,\theta)=(1/2)\|H(\alpha,\theta)+s-\kappa\|^{2}$, i.e., $0\in J_{V}(\alpha^{k},s^{k},\theta^{k})$. We thus converge to a stationary point of the constraint violation function, as desired.
@@ E.1 Conjugate of a general function
For a linear function $g(x,u)=(Pu+a)^{T}x$, we compute the conjugate as
$[-g]^{\star}(x,h)=\sup_{u}h^{T}u-(a+Pu)^{T}x=\begin{dcases}a^{T}x&\text{ if}\;h=-P^{T}x\\ \infty&\text{ otherwise}.\end{dcases}$
@@
For a general function $g(x,u)=\sum_{j=1}^{J}g_{j}(u,x)$, where each $g_{j}$ is concave in $u$, by the theorem of infimal convolutions (Rockafellar and Wets,, 1998, Theorem 11.23(a), p. 493), we compute the conjugate as
$[-g]^{\star}(x,h)=\begin{cases}\underset{h_{j},\dots,h_{J}}{\inf}&\sum_{j=1}^{J}[-g_{j}]^{\star}(x,h_{j})\\ \mbox{subject to}&\sum_{j=1}^{J}h_{j}=h.\end{cases}$
@@
$g(x,u)=\max_{l=1,\dots,L}g_{l}(x,u).$
In practice, these functions $g(x,u)$ are often affine functions. For all the reformulations below, we assume $g(x,u)$ to be of this maximum of concave form.
@@ E.2 Box and Ellipsoidal reformulations
$\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}&g(x,Az+b)\\ \mbox{subject to}&\|z\|_{p}\leq 1\end{array}$
$=\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}~{}\underset{\lambda\geq 0}{\inf}~{}g(x,Az+b)-\lambda\|z\|_{p}+\lambda\\ \end{array}$
@@
$=\begin{array}[t]{ll}\underset{\lambda_{l}\geq 0,\tau}{\inf}&\tau\\ \mbox{subject to}&\lambda_{l}+\underset{Az+b\in S}{\sup}g_{l}(x,Az+b)-\lambda_{l}\|z\|_{p}\leq\tau,\quad l=1,\dots,L,\end{array}$
where $p=\infty$ for box uncertainty and $p\geq 1$ otherwise. We used the Lagrangian in the first equality, and decomposed $g(x,Az+b)$ into its constituents for the second equality. As $l$ and $\lambda$ are separable, in the third step we interchanged the $\inf$ and the $\max$. Next, for each $l$, as the expression is convex in $\lambda$ and concave in $u$, we used the Von Neumann-Fan minimax theorem to interchange the $\inf$ and the $\sup$. We then express the $\max$ in epigraph form by introducing an auxiliary variable $\tau$.
Now, if we define the function $c(z)=\|z\|_{p}$, then using the definition of conjugate functions, we have, for each $l$, the left-hand-side of the constraint equal to
$\lambda_{l}+[-g_{l}(x,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda_{l}c]^{\star}(0).$
@@
$[-g_{l}(x,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda_{l}c]^{\star}(0)=\underset{\omega_{l},h_{l},y_{l}}{\inf}([-g_{l}]^{\star}(x,h_{l})$ $-b^{T}(h_{l}+\omega_{l})+\sigma_{S}(\omega_{l})+\mathcal{I}_{\|y_{l}\|_{q}\leq\lambda_{l}}$
@@
where $h_{l}\in{\mbox{\bf R}}^{m}$, $y_{l}\in{\mbox{\bf R}}^{\hat{m}}$, $\omega_{l}\in{\mbox{\bf R}}^{m}$, and $q$ is the conjugate number of $p$. Substituting these back, we arrive at
$\begin{array}[t]{ll}\underset{x,\tau,h,\omega,y}{\inf}&\tau\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\|y_{l}\|_{q}\leq\tau,\quad l=1,\dots,L\\ &A^{T}(h_{l}+\omega_{l})=y_{l},\quad l=1,\dots,L.\end{array}$
@@
$\begin{array}[t]{ll}\underset{x,h,\omega}{\text{minimize}}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\|A^{T}(h_{l}+\omega_{l})\|_{q}\leq 0,\quad l=1,\dots,L.\end{array}$
@@ E.3 Budget reformulation
$\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}&g(x,Az+b)\\ \mbox{subject to}&\|z\|_{\infty}\leq\rho_{1}\\ &\|z\|_{1}\leq\rho_{2}\\ \end{array}$
@@
$\begin{array}[t]{ll}\underset{\lambda_{l}\geq 0,\tau}{\inf}&\tau\\ \mbox{subject to}&\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+\underset{Az+b\in S}{\sup}g_{l}(x,Az+b)-\lambda_{l1}\|z\|_{\infty}-\lambda_{l2}\|z\|_{1}\leq\tau,\quad l=1,\dots,L.\end{array}$
If we define the functions $c_{1}(z)=\|z\|_{\infty}$, $c_{2}(z)=\|z\|_{1}$, then using the definition of conjugate functions, we have for each $l$ the left-hand-side of the constraint equal to
$\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+[-g_{l}(x,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda_{l1}c_{1}+\lambda_{l2}c_{2}]^{\star}(0)$
$=\underset{h_{l},\omega_{l},y_{l}}{\inf}\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+[-g_{l}]^{\star}(x,h_{l})-b^{T}(h_{l}+\omega_{l})+\sigma_{S}(\omega_{l})+[\lambda_{l1}c_{1}+\lambda_{l2}c_{2}]^{\star}(-y_{l})$
@@
$=\underset{h_{l},\omega_{l},y_{l},r_{l}}{\inf}\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+[-g_{l}]^{\star}(x,h_{l})-b^{T}(h_{l}+\omega_{l})+\sigma_{S}(\omega_{l})+\mathcal{I}_{\|r_{l}-y_{l}\|_{1}\leq\lambda_{l1}}+\mathcal{I}_{\|r_{l}\|_{\infty}\leq\lambda_{l2}}$
@@
$\begin{array}[t]{ll}\underset{x,h_{l},\omega_{l},y_{l},r_{l},\tau}{\inf}&\tau\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\rho_{1}\|r_{l}-y_{l}\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq\tau,\\ &\hfill k=1,\dots,K\\ &A^{T}(h_{l}+\omega_{l})=y_{l},\quad l=1,\dots,L.\end{array}$
@@
$\begin{array}[t]{ll}{\text{minimize}}&f(x)\\ \mbox{subject to}&[-g_{l}]^{\star}(x,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\rho_{1}\|r_{l}-A^{T}(h_{l}+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\\ &\hfill l=1,\dots,L.\end{array}$

### v5

@@ Appendix D Proof of convergence of the augmented Lagrangian algorithm
From step 1, we have a stationary point $(\alpha^{k},\theta^{k})$ of the augmented Lagrangian function at iteration $k$ of Algorithm 1, with fixed multiplier and penalty parameters $\lambda^{k-1}$ and $\mu^{k-1}$. Further properties of this stationary point depend on the convergence, or lack thereof, of the sequence $\{\lambda^{k}\}$. As the step size parameter of $\lambda$ is the penalty $\mu$, we consider the two cases, (a) where the sequence of penalty parameters $\{\mu^{k}\}$ is bounded, and (b) where $\{\mu^{k}\}$ is unbounded. We first analyze case (a). By lines 6-12 of Algorithm 1, we note that for $\{\mu^{k}\}$ to be bounded, the penalty parameter must remain constant from some iteration onwards. By the condition on line 6, with $\tau<1$, this implies $\lim_{k\rightarrow\infty}\|H(\alpha^{k},\theta^{k})\|_{2}=0$, which in turn implies $\lim_{k\rightarrow\infty}\mu^{k}(H(\alpha^{k},\theta^{k}))=0$. This sequence of updates for $\lambda$ is geometric and summable, so the sequence $\{\lambda^{k}\}$ is bounded, with a convergent subsequence. Given the boundedness assumption on the true Lagrangian multiplier $\lambda^{\star}$ in Assumption 3.4, we conclude that the whole sequence converges, ${\it i.e.},$$\{\lambda^{k}\}\rightarrow\lambda^{\star}$. We now turn to case (b), where $\{\mu^{k}\}$ is unbounded. In this case, $\{\lambda^{k}\}$ may not converge. However, by the $\lambda$ update rule in lines 7 of Algorithm 1, we note that while $\{\mu^{k}\}\rightarrow\infty$, $\{\lambda^{k}\}$ is bounded between $\lambda^{\rm{min}}$ and $\lambda^{\rm{max}}$, the safeguard values.
@@
We now show that either the sequence $\{\theta^{k}\}$ converges to a feasible point of the $\widehat{\mathop{\bf CVaR}}$ condition, or $\{\alpha^{k},\theta^{k}\}$ converges to a stationary point of the constraint violation function. We begin with the former. Suppose the sequence of penalty parameters $\{\mu^{k}\}$ is bounded, and converges to $\mu^{\star}$. Then, as $k\to\infty$, $\{\lambda^{k}\}$ converges to the stationary point $\lambda^{\star}$. From step 1, we have that $\{\theta^{t}\}$ in the inner loop converges almost surely to $\theta^{k}=\theta^{\star}(\lambda^{k-1},\mu^{k-1})$ along each step of $\lambda$ and $\mu$’s convergence. In order to establish the convergence of $\{\theta^{k}\}$ in the outer loop to a stationary point $\theta^{\star}(\lambda^{\star},\mu^{\star})$, a feasible point of the $\widehat{\mathop{\bf CVaR}}$ condition (12), we must ensure that $\theta^{\star}(\lambda,\mu)$ is well-behaved, by showing its continuity with respect to $(\lambda,\mu)$.
@@
Recall Assumption 3.3 on the uniqueness of the inner RO solutions $z(\theta,y)$, Assumption 3.4 on the compactness of the set $\mathcal{Z}$, and the convexity (and therefore continuity) of the objective and constraint functions $f$ and $g$ in $x$. Suppose these assumptions hold. By Theorem 4.2.1 of Bank et al., (1984), we establish the continuity of $L(\alpha,\theta,\lambda,\mu)$ in $\theta$, for $\theta\in\mathcal{D}_{\theta}$. As $L$ is bilinear and therefore continuous in $(\lambda,\mu)$, we obtain our desired stability results, i.e., as $\{\lambda^{k}\}$ and $\{\mu^{k}\}$ converge to $\lambda^{\star}$, $\mu^{\star}$, the optimal solution with respect to $\theta$ is continuous, and converges almost surely to the local minimizer $\theta^{\star}$. Note that $\{\alpha^{k}\}$ does not necessarily converge to a stationary point, as the minimum over $\alpha$ is not necessarily unique (Uryasev and Rockafellar,, 2001, Theorem 2). However, by the convergence of the inner loop in step 1 and by the convergence condition of $\mu$ in step 2, as $k\to\infty$ in Algorithm 1, we have a point $(\alpha^{k},\theta^{\star},\lambda^{\star},\mu^{\star})$ satisfying $0\in J_{L}(\alpha^{k},\theta^{\star})$ and $H(\alpha^{k},\theta^{\star})=0$. We can conclude that $\{\theta^{k}\}$ converges almost surely to a feasible point $\theta^{\star}$ of the $\widehat{\mathop{\bf CVaR}}$ condition, where the minimizer over $\alpha$ takes some value $\alpha^{\star}\leq\alpha^{k}$.
@@
On the other hand, suppose the sequence of penalty parameters $\{\mu^{k}\}$ is unbounded. By step 1, for any $\lambda^{k}$ and $\mu^{k}$, we have obtained a stationary point of the augmented Lagrangian function, satisfying $0\in J_{L}(\alpha^{k},\theta^{k})$. Note that, with $\{\mu^{k}\}$ unbounded and $\{\lambda^{k}\}$ bounded, we observe $\lim_{k\rightarrow\infty}\lambda^{k}/\mu^{k}=0$. Therefore, dividing both sides of the conservative Jacobian by $\mu^{k}$ and taking the limit, we obtain:
@@
00 0 $\in\lim_{k\rightarrow\infty}\frac{1}{\eta N}\sum_{x^{i}\in X^{N}}\mathbf{1}\left\{(g(z^{i},u^{i},x^{i})-\alpha^{k})_{+}/\eta+\alpha^{k}-\kappa\geq 0\right\}\mathbf{1}\left\{g(z^{i},u^{i},x^{i})\geq\alpha^{k}\right\}$
$\hskip 85.35826pt((g(z^{i},u^{i},x^{i})-\alpha^{k})_{+}/\eta+\alpha^{k}-\kappa)J_{g}(z^{i})J_{z^{i}}(\theta^{k}),$
@@
00 0 $\in\lim_{k\rightarrow\infty}\frac{1}{N}\sum_{(u^{i},x^{i})\in W^{N}}\frac{1}{\eta N}\sum_{x^{i}\in X^{N}}\mathbf{1}\left\{(g(z^{i},u^{i},x^{i})-\alpha^{k})_{+}/\eta+\alpha^{k}-\kappa\geq 0\right\}$
$\hskip 85.35826pt\left(1-(1/\eta)\mathbf{1}\left\{g(z^{i},u^{i},x^{i})\geq\alpha^{k}\right\}\right)((g(z^{i},u^{i},x^{i})-\alpha^{k})_{+}/\eta+\alpha^{k}-\kappa).$
@@
These are the stationary conditions for the constraint violation function $V(\alpha,\theta)=(1/2)\|H(\alpha,\theta)\|^{2}$, i.e., $0\in J_{V}(\alpha^{k},\theta^{k})$. We thus converge to a stationary point of the constraint violation function, as desired.
@@ E.1 Conjugate of a general function
For a linear function $g(z,u)=(Pu+a)^{T}z$, we compute the conjugate as
$[-g]^{\star}(z,h)=\sup_{u}h^{T}u-(a+Pu)^{T}z=\begin{dcases}a^{T}z&\text{ if}\;h=-P^{T}z\\ \infty&\text{ otherwise}.\end{dcases}$
@@
For a general function $g(z,u)=\sum_{j=1}^{J}g_{j}(u,z)$, where each $g_{j}$ is concave in $u$, by the theorem of infimal convolutions (Rockafellar and Wets,, 1998, Theorem 11.23(a), p. 493), we compute the conjugate as
$[-g]^{\star}(z,h)=\begin{cases}\underset{h_{j},\dots,h_{J}}{\inf}&\sum_{j=1}^{J}[-g_{j}]^{\star}(z,h_{j})\\ \mbox{subject to}&\sum_{j=1}^{J}h_{j}=h.\end{cases}$
@@
$g(z,u)=\max_{l=1,\dots,L}g_{l}(z,u).$
In practice, these functions $g(z,u)$ are often affine functions. For all the reformulations below, we assume $g(z,u)$ to be of this maximum of concave form.
@@ E.2 Box and Ellipsoidal reformulations
$\begin{array}[t]{ll}\underset{Av+b\in S}{\sup}&g(z,Av+b)\\ \mbox{subject to}&\|v\|_{p}\leq 1\end{array}$
$=\begin{array}[t]{ll}\underset{Av+b\in S}{\sup}~{}\underset{\lambda\geq 0}{\inf}~{}g(z,Av+b)-\lambda\|v\|_{p}+\lambda\\ \end{array}$
@@
$=\begin{array}[t]{ll}\underset{\lambda_{l}\geq 0,\tau}{\inf}&\tau\\ \mbox{subject to}&\lambda_{l}+\underset{Av+b\in S}{\sup}g_{l}(z,Av+b)-\lambda_{l}\|v\|_{p}\leq\tau,\quad l=1,\dots,L,\end{array}$
where $p=\infty$ for box uncertainty and $p\geq 1$ otherwise. We used the Lagrangian in the first equality, and decomposed $g(z,Av+b)$ into its constituents for the second equality. As $l$ and $\lambda$ are separable, in the third step we interchanged the $\inf$ and the $\max$. Next, for each $l$, as the expression is convex in $\lambda$ and concave in $u$, we used the Von Neumann-Fan minimax theorem to interchange the $\inf$ and the $\sup$. We then express the $\max$ in epigraph form by introducing an auxiliary variable $\tau$.
Now, if we define the function $c(v)=\|v\|_{p}$, then using the definition of conjugate functions, we have, for each $l$, the left-hand-side of the constraint equal to
$\lambda_{l}+[-g_{l}(z,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda_{l}c]^{\star}(0).$
@@
$[-g_{l}(z,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda_{l}c]^{\star}(0)=\underset{\omega_{l},h_{l},y_{l}}{\inf}([-g_{l}]^{\star}(z,h_{l})$ $-b^{T}(h_{l}+\omega_{l})+\sigma_{S}(\omega_{l})+\mathcal{I}_{\|y_{l}\|_{q}\leq\lambda_{l}}$
@@
where $h_{l}\in{\mathbf{R}}^{m}$, $y_{l}\in{\mathbf{R}}^{\hat{m}}$, $\omega_{l}\in{\mathbf{R}}^{m}$, and $q$ is the conjugate number of $p$. Substituting these back, we arrive at
$\begin{array}[t]{ll}\underset{z,\tau,h,\omega,y}{\inf}&\tau\\ \mbox{subject to}&[-g_{l}]^{\star}(z,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\|y_{l}\|_{q}\leq\tau,\quad l=1,\dots,L\\ &A^{T}(h_{l}+\omega_{l})=y_{l},\quad l=1,\dots,L.\end{array}$
@@
$\begin{array}[t]{ll}\underset{z,h,\omega}{\text{minimize}}&f(z)\\ \mbox{subject to}&[-g_{l}]^{\star}(z,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\|A^{T}(h_{l}+\omega_{l})\|_{q}\leq 0,\quad l=1,\dots,L.\end{array}$
@@ E.3 Budget reformulation
$\begin{array}[t]{ll}\underset{Az+b\in S}{\sup}&g(z,Av+b)\\ \mbox{subject to}&\|v\|_{\infty}\leq\rho_{1}\\ &\|v\|_{1}\leq\rho_{2}\\ \end{array}$
@@
$\begin{array}[t]{ll}\underset{\lambda_{l}\geq 0,\tau}{\inf}&\tau\\ \mbox{subject to}&\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+\underset{Av+b\in S}{\sup}g_{l}(z,Av+b)-\lambda_{l1}\|v\|_{\infty}-\lambda_{l2}\|v\|_{1}\leq\tau,\quad l=1,\dots,L.\end{array}$
If we define the functions $c_{1}(v)=\|v\|_{\infty}$, $c_{2}(v)=\|v\|_{1}$, then using the definition of conjugate functions, we have for each $l$ the left-hand-side of the constraint equal to
$\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+[-g_{l}(z,A\cdot+b)+\chi_{S}(A\cdot+b)+\lambda_{l1}c_{1}+\lambda_{l2}c_{2}]^{\star}(0)$
$=\underset{h_{l},\omega_{l},y_{l}}{\inf}\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+[-g_{l}]^{\star}(z,h_{l})-b^{T}(h_{l}+\omega_{l})+\sigma_{S}(\omega_{l})+[\lambda_{l1}c_{1}+\lambda_{l2}c_{2}]^{\star}(-y_{l})$
@@
$=\underset{h_{l},\omega_{l},y_{l},r_{l}}{\inf}\lambda_{l1}\rho_{1}+\lambda_{l2}\rho_{2}+[-g_{l}]^{\star}(z,h_{l})-b^{T}(h_{l}+\omega_{l})+\sigma_{S}(\omega_{l})+\mathcal{I}_{\|r_{l}-y_{l}\|_{1}\leq\lambda_{l1}}+\mathcal{I}_{\|r_{l}\|_{\infty}\leq\lambda_{l2}}$
@@
$\begin{array}[t]{ll}\underset{z,h_{l},\omega_{l},y_{l},r_{l},\tau}{\inf}&\tau\\ \mbox{subject to}&[-g_{l}]^{\star}(z,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\rho_{1}\|r_{l}-y_{l}\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq\tau,\\ &\hfill k=1,\dots,K\\ &A^{T}(h_{l}+\omega_{l})=y_{l},\quad l=1,\dots,L.\end{array}$
@@
$\begin{array}[t]{ll}{\text{minimize}}&f(z)\\ \mbox{subject to}&[-g_{l}]^{\star}(z,h_{l})+\sigma_{S}(\omega_{l})-b^{T}(h_{l}+\omega_{l})+\rho_{1}\|r_{l}-A^{T}(h_{l}+\omega_{l})\|_{1}+\rho_{2}\|r_{l}\|_{\infty}\leq 0,\\ &\hfill l=1,\dots,L.\end{array}$